# Exploring Spatial Segregation Induced by Competition Avoidance as Driving Mechanism for Emergent Coexistence in Microbial Communities

### 12 Feb 2024 v2

@@ Abstract
This study investigates the role of spatial segregation, prompted by competition avoidance, as a key mechanism for emergent coexistence within microbial communities. Recognizing these communities as complex adaptive systems, we challenge the sufficiency of pairwise interaction models and consider the impact of spatial dynamics. We developed an individual-based spatial simulation depicting bacterial movement through a pattern of random walks influenced by competition avoidance, leading to the formation of spatially segregated clusters. This model was integrated with a Lotka-Volterra metapopulation framework focused on competitive interactions. Our findings reveal that spatial segregation alone can lead to emergent coexistence in microbial communities, offering a new perspective on the formation of stable, coexisting microbe clusters that differ significantly from their behavior in isolated pairwise interactions. This study underscores the importance of considering spatial factors in understanding the dynamics of microbial ecosystems.
@@ I Introduction
Following this direction, we developed an individual-based spatial simulation to depict the individual movement of bacteria, leading to the formation of spatially segregated clusters resulting from the escape from regions with high competition. This simulation was then utilized to calculate the initial conditions for a metapopulations Lotka-Volterra model with only competitive interactions, capturing the growth dynamics of such patches of bacteria. This study shows that: i) segregation of clusters of bacteria can be obtained as a result of competition avoidance only and, therefore, it can potentially occur in any conditions regardless of the environmental setup; ii) building upon the latter justification, a metapopulation Lotka-Volterra formalism can be adopted and this alters considerably the pattern of coexistence for the species at the equilibrium. Specifically, for low diffusion rates and high compositional heterogeneity among patches, it is possible to reproduce the emergence observed by Chang et al., i.e. the formation of stable communities formed of by multiple microbial species, the majority of which do not coexist when isolated in pairwise combinations. As a benchmark to validate our model, we also confirmed its ability to replicate the three macroecological laws governing microbial communities, as discovered by J. Grilli[23].
@@ II Individual-based Spatial Simulation for Bacterial Motion
In other words we model bacterial movement selecting a random direction and an intensity proportional to the level of competitiveness, i.e. the number of competitors $N_{c}^{i}$ in the neighborhood of the node $i$. Taking inspiration from neural networks, the intensity of motion is modelled as a sigmoid function (figure 1). The parameter $\alpha$ controls the the shape of the curve; if $\alpha\to\infty$ the function tends to the Heaviside step-function. The threshold parameter $N_{th}$ controls the position of the inflection point $N_{c}^{i}=N_{th}$. This function cannot be equal to zero, disallowing bacteria to be completely still, and its maximum value is equal to $R$, i.e. the radius of the neighborhood area.
Figure 1: The function for the bacterial intensity of motion.
Each curve is described by equation 2 with different $\alpha$ between 1 and 20. The greater $\alpha$, the more the function tends to a step function. Here, $R$ is chosen equal to 0.2.
The simulation stops when all the bacteria minimize their intensity motion, i.e. when their activity results in a random walk confined in a small portion of space. As we will see later, it is always possible to choose the parameters $(\alpha,N_{th})$ to ensure that this formalism naturally generates spatially segregated clusters of bacteria, whose dimensions depends on the radius of interactions $R$.
In order to quantitatively compute the different clusters of bacteria in the final network we employed the classical Louvain Algorithm for communities detection [25]. It optimizes a quality function known as modularity, which quantifies the strength of the community structure in a network. The algorithm optimizes modularity by iteratively moving nodes between communities, enhancing the overall cohesion within communities while reducing connectivity between them. The Louvain algorithm is a popular choice in the field of network science, despite the numerous limitations associated with modularity optimization algorithms [26]. In our context, the selection of a particular algorithm has negligible effects.
@@ III.1 On the stability of Lotka-Volterra equations
Mathematical analyses of stability typically focus on conditions near equilibrium points due to analytical challenges in dealing with nonlinear systems at a distance from equilibrium. For this reason, Holling [35] suggested defining the behavior of ecological systems with two distinct properties: resilience and stability. Resilience pertains to the persistence of relationships within a system and measures its ability to absorb changes in state variables, driving variables, and parameters while still persisting. Stability, on the other hand, refers to a system’s capacity to return to equilibrium after a temporary disturbance, with a more rapid and less fluctuating return indicating greater stability. The well-known mutual invasion criterion, often associated with the concept of resilience, serves as a notable benchmark. For stable coexistence, it demands that each species within a community demonstrates positive population growth rates when invading a pre-existing community of competitors from low density [36]. This is exactly the indicator used by Chang et al. to assess stable coexistence in their experimental communities of microbial species. Take a moment to focus on a scenario involving only two species within the system. In the context of competitive Lotka-Volterra equations, it is established that the invasion criterion holds true when $A_{12}<K_{1}/K_{2}$ and $A_{21}<K_{2}/K_{1}$ [37]. Consequently, note that weak interactions with zero mean would almost always satisfy the invasion criterion and, consequently, lead to stable coexistence. However, Chang et al. [6] observed in their experiments that only a relatively small fraction (about 30%) of possible pairs of species complies with the invasion criterion. When contemplating the parameters $A_{ij}$ as ‘universal’ coefficients for pairwise interactions, it appears that these experimental results are scarcely consistent with the conditions necessary for mathematical asymptotic stability. The work by Abramson and Zanette [38], provides us a workaround. They randomly selected interaction coefficients for a system comprising $N$ Lotka-Volterra species from a uniform distribution centered around one. Remarkably, the resulting phase space exhibited a multitude of fixed points, with a majority featuring both positive and negative eigenvalues—indicating instability. Consequently, the system traverses various unstable equilibria, leading to instances where the population of certain species undergoes pseudo-extinctions, reaching very low concentrations before rebounding. The noteworthy finding in their study is the demonstration that introducing a lower bound, $X_{0}$, to the populations induces a shift in the stability of nearly all equilibria, transforming them into stable states. It is important to note that the imposition of a lower threshold on populations is entirely justified. Indeed, the population density of a species or genotype confined to a specific spatial volume $V$ cannot fall below $V^{-1}$ unless it disappears entirely. When describing densities, it becomes essential to establish a threshold below which the density effectively approaches zero. They additionally demonstrated that enlarging the domain from which interaction coefficients are sampled results in a reduction in the number of species surviving at equilibrium. All of these heuristic discussions provide compelling justifications for our parameter choices in the model, which will be detailed in the results section. Moreover, we recall that our theoretical setup incorporates the structural aspect of patches, a factor that significantly influences coexistence and stability requirements. The next section will provide further clarification on this point.
@@ III.2 Mesoscopic Interpretation of the Interactions Parameters
where we have considered $A_{ii}=1$. To comply with the mutual invasion criterion the same must hold also for $A_{21}$, inverting the indices. However species 2, in absence of species 1, has grown equally in all the patches until equilibrium; this means that $\phi^{*}_{2\alpha}=1/N_{p},\,\,\forall\alpha$. Considering also that $\sum_{\alpha}\phi_{1\alpha}=1$, we obtain that the factor $\Omega$ is equal to one and we revert to the classical invasion criterion conditions. Therefore, this elucidates how the mesoscale structure can at the same time relevantly affect the global interactions $A_{ij}^{G}$, and so the steady state, for species in communities while maintaining the same conditions for pairwise stable coexistence in the sense of the invasion criterion.
@@ IV Results
Figure 2: Emergence of Spatially Segregated Bacterial Clusters Resulting from Competition Avoidance.
@@
The simulation in panel a) was conducted with $R=0.2$ and $N_{th}=60$. We considered $n=2000$ nodes randomly assigned to $N=100$ different species, each one distringuished by a different color. The top row presents the random geometric network representation of the system, while the bottom row illustrates the intensity motion distribution. In the smaller panel b), the graph at different time steps is shown for $R=0.3$ and $N_{th}=160$. The last two snapshots in b) and panel c) show the application of the Louvain algorithm to the RGG at the equilibrium, identifying the different patches. The dimension of the patches is proportional to the number of nodes within them and the widths of the edges are chosen accordingly to the number of links between the respective communities. All simulations are performed in a two-dimensional square of dimensions $2\times 2$.
@@
Upon reaching equilibrium, each simulation yields the initial distribution of various species across patches, representing the initial conditions in our metapopulation model for species growth. We numerically solved the system of differential equations using an explicit Runge-Kutta method of order 5. For simplicity, we opted for identical and unitary intrinsic growth rates and carrying capacities for all species ($K_{i}=1,\,\,r_{i}=1,\,\,\forall i$). In light of section III.A, we chose not to require asymptotic stability but to follow a direction similar to Abramson et al. [38], therefore we randomly sampled the off-diagonal entries of the interaction matrix from a lognormal distribution with a mean ($\mu$) set to one, varying its variance ($\sigma^{2}$) from 0.05 to 0.5 ($A_{ij}\sim\text{Lognormal}(\mu,\sigma)$). The lognormal distribution was chosen in order to sample only positive $A_{ij}$. The diagonal elements were uniformly set to one ($A_{ii}=1$). Additionally, we incorporated a lower limit for the population, denoted as $X^{0}_{i\alpha}$, set at $10^{-3}$. Furthermore, we explored various values for the diffusion parameter ($\mu$), spanning from zero to one.
Figure 3: Graphical Description of the Simulation Process.
The paper aims to replicate the experimental procedure outlined by Chang et al. Consequently, after conducting the spatial simulation, we utilize the metapopulation Lotka-Volterra formalism to model species growth. The initial plot on the left illustrates the temporal dynamics of relative abundances for each species. Following each simulation, we identify the species that survive at equilibrium and subsequently consider all potential pairs of these species, re-executing the entire model (including spatial simulation and growth) using only those pairs. Mirroring the experimental setup, we explore three distinct initial distribution proportions for the pairs (50%-50%, 95%-5%, 5%-95%) to determine if they can coexist under each scenario. The plots on the right depict the spatial simulations for the two species (top) and their growth dynamics across the three distributions (bottom).
Confirming the findings of Abramson et al. [38], in this setup, we observed that the numerical simulations display equilibrium. We considered the system to exhibit equilibrium dynamics if the coefficient of variation for the populations of the species was below 1% in the last 500 time-steps. This criterion was met in the majority of cases, while in some instances, species’ populations exhibited oscillatory behavior. We focused solely on the former cases for our analysis. Figure 3 provides a visual summary of the simulation process’s structure. The idea is to replicate the procedure in [6]: after completing the spatial simulation, we employ the metapopulation Lotka-Volterra formalism to simulate species growth. Subsequently, we identify the species that reach equilibrium and proceed to re-execute the entire model again, including both spatial simulation and growth, considering all the possible pairs of survivors. Consistent with the experimental design, we investigate three different initial distribution proportions for the pairs (50%-50%, 95%-5%, 5%-95%) to assess their potential for coexistence under each scenario in the absence of the other species.
The initial series of simulations consisted of 100 runs, considering different combinations of the interaction standard deviation ($\sigma$) along with the diffusion parameter ($\mu$). We examined three values of $\sigma$ (0.05, 0.1, and 0.5) and three values of $\mu$ (0.1, 0.3, and 0.5). Figure 4 displays the outcomes for the number of survivors while varying $\sigma$ and keeping $\mu$ constant (top row), and vice versa (bottom row). A vertical dashed red line indicates the mean number of surviving species over 100 runs of the classical Mean-Field Lotka-Volterra without patch structure. Notably, across all cases, the patch structure consistently facilitates a significantly higher number of surviving species compared to the mean field scenario, indicative of enhanced coexistence among species. The distribution shifts towards higher values for the number of survivors with higher $\sigma$ and lower diffusion rates $\mu$. This latter trend can be explained straightforwardly: as $\mu$ increases, species can explore the entire patch network more rapidly, converging towards the classical mean-field approach. Conversely, when variance increases, the distribution exhibits heavier tails, and the correlation between higher variances and a greater number of survivors suggests that increased heterogeneity in the interaction patterns among species, and therefore in the composition of the patches, can promote coexistence. We will revisit the effect of the heterogeneity later in our analysis.
Figure 4: Distributions of the Number of Survivors.
[…]
@@
Further insights can be obtained looking to the fractions $\phi_{i}$ before and after executing the model. Specifically, refer to figure 5 to observe the ‘proximity measure’ $S_{ij}$, as defined in equation 5, for each pair of species both before and after allowing them to grow and migrate. In the initial state, $S_{ij}$ exhibits a bell-shaped distribution, indicating that the spatial simulation results in species that are, on average, evenly segregated among the patches. Following the model execution, the surviving species demonstrate a distribution more concentrated around zero and with a long tail, suggesting that survivors are more likely to be highly segregated with a small minority of them being able to coexist in the same patches. This phenomenon facilitates the coexistence of species, even when their net interaction $A_{ij}$ is high.
Figure 5: Distribution for the Proximity Measure.
@@
Distribution of the proximity measure $S_{ij}$, as defined in equation 5 for each possible pair of species, before and after species’ growth and over 100 runs of the complete model. In the final state we considered all possible couples of only survived species.
@@
Figure 6: Number of Survivors varying the Initial Distribution of the Species and the Diffusion Rate.
The heatmap depicts the number of survivors at the equilibrium varying the initial distribution of the species, through the concentration parameter $a$ in a Dirichlet distribution (y-axis in log scale), and the diffusion rate $\mu$ (x-axis). Brighter colors indicate higher values for the number of survivors. The white horizontal dashed line indicates the value of $a$ which best fits with the initial distribution of the species provided by our spatial simulation. For each pair of values $(a,\mu)$ we performed 100 runs of the model with $K_{i}=1$ and $r_{i}=1\,\forall i$, $X^{0}_{i\alpha}=10^{-3}\,\forall\,(i,\,\alpha)$.
@@
Figure 7: Distribution for the Percentage of Coexisting Pairs of Species.
Each box represents the distribution for the percentage of coexisting species’ pairs over 100 runs for different combinations of $\mu$ and $\sigma$, as depicted along the x-axis. The box extends from the first quartile (Q1) to the third quartile (Q3) of the data, with a line at the median. The whiskers extend from the box to the farthest data point lying within 1.5x the inter-quartile range (IQR) from the box. Flier points indicate outliers. The red horizontal dashed line indicates the experimental value for the percentage of coexisting pairs found in [6].
In fact, in the absence of multiple species, spatial self-organization in this scenario leads to a random process distributing both species equally across all patches with high probability (see picture on the left of figure 3). This implies an unaffected interaction pattern, rendering the spatial distribution incapable of promoting coexistence. When there is a significant disparity in the initial concentration of the two species, as previously mentioned, the patch structure does not impact the invasion criterion when one species invades the other. Consequently, the mesoscale structure, while potentially influencing coexistence in multi-species configurations, does not change the criteria for coexistence in the case of two species. This elucidates the findings illustrated in Figure 7.
@@
Figure 8: Macroecological laws.
We conducted multiple model runs, sampling different interactions $A_{ij}$ in each simulation. A total of 100 runs were performed for three distinct values of $\mu$ (0.1, 0.3, and 0.5) in the top row, and for three different values of $\sigma$ (0.05, 0.1, 0.5) in the bottom one. Subsequent to all the simulations, we assessed whether the species’ abundances across runs followed a gamma distribution (left figure), if the mean and variance of the abundances were related through Taylor’s law with an exponent of two (central figure), and finally, if the mean abundance was distributed according to a lognormal distribution (right figure). The analysis resulted in mean $R^{2}$ values of 0.97, 0.5, and 0.86, respectively, for the three laws. The various colors of the circles represent the three values of $\mu$ in the top row and of $\sigma$ in the bottom one. The remaining parameters in the model were selected as in Figure 4.

### 26 Feb 2024 v3

@@ Abstract
This study investigates the role of spatial segregation, prompted by competition avoidance, as a key mechanism for emergent coexistence within microbial communities. Recognizing these communities as complex adaptive systems, we challenge the sufficiency of mean-field pairwise interaction models and consider the impact of spatial dynamics. We developed an individual-based spatial simulation depicting bacterial movement through a pattern of random walks influenced by competition avoidance, leading to the formation of spatially segregated clusters. This model was integrated with a Lotka-Volterra metapopulation framework focused on competitive interactions. Our findings reveal that spatial segregation combined with low diffusion rates and high compositional heterogeneity among patches can lead to emergent coexistence in microbial communities. This reveals a novel mechanism underpinning the formation of stable, coexisting microbe clusters, which is nonetheless incapable of promoting coexistence in the case of isolated pairs of species. This study underscores the importance of considering spatial factors in understanding the dynamics of microbial ecosystems.
@@ I Introduction
Following this direction, we developed an individual-based spatial simulation to depict the individual movement of bacteria, leading to the formation of spatially segregated clusters resulting from the escape from regions with high competition. Under the assumption of time-scales separation between movement and growth, the result of the spatial simulation was then utilized to calculate the initial conditions for a metapopulations Lotka-Volterra model with only competitive interactions, capturing the growth dynamics of such patches of bacteria. This study shows that: i) segregation of clusters of bacteria can be obtained as a result of competition avoidance only and, therefore, it can potentially occur in any conditions regardless of the environmental setup; ii) building upon the latter justification, a metapopulation Lotka-Volterra formalism can be adopted and this alters considerably the pattern of coexistence for the species at the equilibrium. Specifically, for low diffusion rates and high compositional heterogeneity among patches, it is possible to reproduce the emergence observed by Chang et al., i.e. the formation of stable communities formed by multiple microbial species, the majority of which do not coexist when isolated in pairwise combinations. As a benchmark to validate our model, we also confirmed its ability to replicate the three macroecological laws governing microbial communities, as discovered by J. Grilli[23].
@@ II Individual-based Spatial Simulation for Bacterial Motion
Stated differently, we model bacterial movement selecting a random direction and an intensity proportional to the level of competitiveness, i.e. the number of competitors $N_{c}^{i}$ in the neighborhood of the node $i$. Taking inspiration from neural networks, the intensity of motion is modelled as a sigmoid function. The parameter $\alpha$ controls the the shape of the curve; if $\alpha\to\infty$ the function tends to the Heaviside step-function. The threshold parameter $N_{th}$ controls the position of the inflection point $N_{c}^{i}=N_{th}$. This function cannot be equal to zero, disallowing bacteria to be completely still, and its maximum value is equal to $R$, i.e. the radius of the neighborhood area. The simulation stops when all the bacteria minimize their intensity motion, i.e. when their activity results in a random walk confined in a small portion of space. In order to quantitatively compute the different clusters of bacteria in the final network we employed the classical Louvain Algorithm for communities detection [25]. It optimizes a quality function known as modularity, which quantifies the strength of the community structure in a network. The algorithm optimizes modularity by iteratively moving nodes between communities, enhancing the overall cohesion within communities while reducing connectivity between them. The Louvain algorithm is a popular choice in the field of network science, despite the numerous limitations associated with modularity optimization algorithms [26]. In our context, the selection of a particular algorithm has negligible effects.
@@ III.1 On the stability of Lotka-Volterra equations
Mathematical analyses of stability typically focus on conditions near equilibrium points due to analytical challenges in dealing with nonlinear systems at a distance from equilibrium. For this reason, Holling [35] suggested defining the behavior of ecological systems with two distinct properties: resilience and stability. Resilience pertains to the persistence of relationships within a system and measures its ability to absorb changes in state variables, driving variables, and parameters while still persisting. Stability, on the other hand, refers to a system’s capacity to return to equilibrium after a temporary disturbance, with a more rapid and less fluctuating return indicating greater stability. The well-known mutual invasion criterion, often associated with the concept of resilience, serves as a notable benchmark. For stable coexistence, it demands that each species within a community demonstrates positive population growth rates when invading a pre-existing community of competitors from low density [36]. This is exactly the indicator used by Chang et al. to assess stable coexistence in their experimental communities of microbial species. Take a moment to focus on a scenario involving only two species within the system. In the context of competitive Lotka-Volterra equations, it is established that the invasion criterion holds true when $A_{12}<K_{1}/K_{2}$ and $A_{21}<K_{2}/K_{1}$ [37]. Consequently, note that weak interactions with zero mean would almost always satisfy the invasion criterion and, consequently, lead to stable coexistence. However, Chang et al. [6] observed in their experiments that only a relatively small fraction (about 30%) of possible pairs of species complies with the invasion criterion. When contemplating the parameters $A_{ij}$ as ‘universal’ coefficients for pairwise interactions, i.e. only depending on species’ types, it appears that these experimental results are scarcely consistent with the conditions necessary for mathematical asymptotic stability. The work by Abramson and Zanette [38], provides us a workaround. They randomly selected interaction coefficients for a system comprising $N$ Lotka-Volterra species from a uniform distribution centered around one. Remarkably, the resulting phase space exhibited a multitude of fixed points, with a majority featuring both positive and negative eigenvalues—indicating instability. Consequently, the system traverses various unstable equilibria, leading to instances where the population of certain species undergoes pseudo-extinctions, reaching very low concentrations before rebounding. The noteworthy finding in their study is the demonstration that introducing a lower bound, $X_{0}$, to the populations induces a shift in the stability of nearly all equilibria, transforming them into stable states. It is important to note that the imposition of a lower threshold on populations is entirely justified. Indeed, the population density of a species or genotype confined to a specific spatial volume $V$ cannot fall below $V^{-1}$ unless it disappears entirely. When describing densities, it becomes essential to establish a threshold below which the density effectively approaches zero. All of these heuristic discussions provide compelling justifications for our parameter choices in the model, which will be detailed in the results section. Moreover, we recall that our theoretical setup incorporates the structural aspect of patches, a factor that significantly influences coexistence and stability requirements. The next section will provide further clarification on this point.
@@ III.2 Mesoscopic Interpretation of the Interactions Parameters
where we have considered $A_{ii}=1$. To comply with the mutual invasion criterion the same must hold also for $A_{21}$, inverting the indices. However species 2, in absence of species 1, has grown equally in all the patches until equilibrium; this means that $\phi^{*}_{2\alpha}=1/N_{p},\,\,\forall\alpha$. Considering also that $\sum_{\alpha}\phi_{1\alpha}=1$, we obtain that the factor $\Omega$ is equal to one and we revert to the classical invasion criterion conditions. This elucidates how the mesoscale structure can at the same time relevantly affect the global interactions $A_{ij}^{G}$, and so the steady state, for species in communities while maintaining the same conditions for pairwise stable coexistence in the sense of the invasion criterion.
@@ IV Results
Figure 1: In this example the individual-based simulation was conducted in a two-dimensional square of dimensions $2\times 2$ with $R=0.2$, $\alpha=7$ and $N_{th}=60$. We considered $n=2000$ nodes (bacteria) randomly assigned to $N=100$ different species. The top row of panel A) presents the random geometric network representation of the system where each species is indicated with a different color, while the bottom row illustrates the intensity motion distribution as described by equation 2, both showed at three different time-steps ($t=0$, $t=t_{end}/2$ and at the end of the simulation $t_{end}$). In panel B), we present the outcome of the community detection algorithm. Initially, we depict the identified communities using distinct colors (top row). Subsequently, we construct a patch network wherein communities are represented by individual nodes of corresponding colors (bottom row). In this network the size of each node is proportional to the number of bacteria within the community and the edges are weighted based on the number of links connecting the communities.
Figure 1 presents an example of simulation with $R=0.2$, $\alpha=7$ and $N_{th}=60$, with snapshots of the system at different time steps. In panel A) The RGG is displayed alongside the associated intensity motion distribution. As the nodes cluster, the distribution shifts towards very low values. This phenomenon persists across all the higher values of $R$ investigated, extending up to 0.5. The straightforward outcome of increasing the distance in the RGG is a reduction in the number of patches and an increase in their dimensions. Consequently, for the subsequent analysis, we focus solely on the case where $R=0.2$. All the simulations are performed in $2\times 2$ bidimensional square. It’s important to note that the range of velocities considered is biologically plausible: bacteria can move at a wide range of speeds ranging from $1\,\,\mu m/s$ to $1000\,\,\mu m/s$ [41]; considering the square as $2\times 2\,\,mm^{2}$, this implies that bacteria at the beginning of our simulation travel on average 141 $\mu m$ in a few seconds of straight-line motion, with a maximum possible value of about 280 $\mu m$; then, at the end of simulation, they all reach velocities of very few $\mu m$ per second.
Upon reaching equilibrium, each simulation yields the initial distribution of various species across patches, representing the initial conditions in our metapopulation model for species growth. We numerically solved the system of differential equations using an explicit Runge-Kutta method of order 5. For simplicity, we opted for identical and unitary intrinsic growth rates and carrying capacities for all species ($K_{i}=1,\,\,r_{i}=1,\,\,\forall i$). In light of section III.A, we chose not to require asymptotic stability but to follow a direction similar to Abramson et al. [38], therefore we randomly sampled the off-diagonal entries of the interaction matrix from a lognormal distribution with a mean ($\mu$) set to one, varying its variance ($\sigma^{2}$) from 0.05 to 0.5 ($A_{ij}\sim\text{Lognormal}(\mu,\sigma)$). The diagonal elements were uniformly set to one ($A_{ii}=1$). Additionally, we incorporated a lower limit for the population, denoted as $X^{0}_{i\alpha}$, set at $10^{-3}$. Furthermore, we explored various values for the diffusion parameter ($\mu$), spanning from zero to one. Confirming the findings of Abramson et al. [38], in this setup, we observed that the numerical simulations display equilibrium. We considered the system to exhibit equilibrium dynamics if the coefficient of variation for the populations of the species was below 1% in the last 500 time-steps. This criterion was met in the majority of cases, while in some instances, species’ populations exhibited oscillatory behavior. We focused solely on the former cases for our analysis. To assess the stability of these equilibrium configurations, we employ the mutual invasion criterion. This involves systematically reducing the abundance of each surviving species at equilibrium one by one and observing whether they exhibit a positive growth rate, ultimately returning to values close to their previous abundance levels. We observed that approximately 90% of the surviving species consistently demonstrate regrowth potential in all the cases investigated, guaranteeing a form of stability for our simulated systems. Those species that became extinct were primarily those with already minimal abundance levels at equilibrium.
The idea is to organize the simulations to replicate the procedure in [6]: the spatial simulation ends up with the RGG at equilibrium, on which we employed the modularity optimization Louvain algorithm which successfully identifies distinct communities, i.e. the different patches of the metapopulation network (panel B of figure 1). Afterwards, we use the metapopulation Lotka-Volterra formalism to simulate species growth and to identify the species that reach equilibrium. Subsequently, we proceed to re-execute the entire model again, including both spatial simulation and growth, considering all the possible pairs of survivors. Consistent with the experimental design, we investigate three different initial distribution proportions for the pairs (50%-50%, 95%-5%, 5%-95%) to assess their potential for coexistence under each scenario in the absence of the other species.
The initial series of simulations consisted of 100 runs, considering different combinations of the interaction standard deviation ($\sigma$) along with the diffusion parameter ($\mu$). We examined three values of $\sigma$ (0.05, 0.1, and 0.5) and three values of $\mu$ (0.1, 0.3, and 0.5). Figure 2 displays the outcomes for the number of survivors ($N_{S}$) while varying $\sigma$ and keeping $\mu$ constant (top row), and vice versa (bottom row). A vertical dashed red line indicates the mean number of surviving species over 100 runs of the classical Mean-Field Lotka-Volterra without patch structure. Notably, across all cases, the patch structure consistently facilitates a significantly higher number of surviving species compared to the mean field scenario, indicative of enhanced coexistence among species. The distribution shifts towards higher values for the number of survivors with higher $\sigma$ and lower diffusion rates $\mu$. This latter trend can be explained straightforwardly: as $\mu$ increases, species can explore the entire patch network more rapidly, converging towards the classical mean-field approach. Conversely, when variance increases, the distribution exhibits heavier tails, and the correlation between higher variances and a greater number of survivors suggests that increased heterogeneity in the interaction patterns among species, and therefore in the composition of the patches, can promote coexistence. We will revisit the effect of the heterogeneity later in our analysis.
@@
Further insights can be obtained looking to the fractions $\phi_{i}$ before and after executing the model. Specifically, refer to figure 3 to observe the ‘proximity measure’ $S_{ij}$, as defined in equation 5, for each pair of species both before and after allowing them to grow and migrate. In the initial state, $S_{ij}$ exhibits a bell-shaped distribution, indicating that the spatial simulation results in species that are, on average, evenly segregated among the patches. Following the model execution, the surviving species demonstrate a distribution more concentrated around zero and with a long tail, suggesting that survivors are more likely to be highly segregated with a small minority of them being able to coexist in the same patches. As described by equation 5, this phenomenon facilitates the coexistence of species, even when their net interaction $A_{ij}$ is high.
@@
Figure 3: Distribution of the proximity measure $S_{ij}$, as defined in equation 5, for each possible pair of species, before and after species’ growth and over 100 runs of the complete model. Given that $S_{ij}$ is not defined when one or both species are extinct, in the final state we considered all possible couples of only survived species. For the simulation here $n=2000$, $N=100$, $R=0.2$, $N_{th}=60$, $\alpha=7$, $\sigma=0.1$, $\mu=0.3$, $K_{i}=1$ and $r_{i}=1\,\forall i$, $X^{0}_{i\alpha}=10^{-3}\,\forall\,(i,\,\alpha)$.
@@
Figure 4: The heatmap depicts the number of survivors ($N_{S}$) at the equilibrium varying the initial distribution of the species, through the concentration parameter $a$ in a Dirichlet distribution (y-axis in log scale), and the diffusion rate $\mu$ (x-axis). Brighter colors indicate higher values for the number of survivors. The white horizontal dashed line indicates the value of $a$ which best fits with the initial distribution of the species provided by our spatial simulation. For each pair of values $(a,\mu)$ we performed 100 runs of the model with $\sigma=0.1$, $K_{i}=1$ and $r_{i}=1\,\forall i$, $X^{0}_{i\alpha}=10^{-3}\,\forall\,(i,\,\alpha)$. To ensure the figure appears continuous, we used quadratic interpolation.
@@
Figure 5: Each box represents the distribution for the percentage of coexisting species’ pairs (picked from the set of survivors in the multi-species simulation) over 100 runs for different combinations of $\mu$ and $\sigma$, as depicted along the x-axis. The box extends from the first quartile (Q1) to the third quartile (Q3) of the data, with a line at the median. The whiskers extend from the box to the farthest data point lying within 1.5x the inter-quartile range (IQR) from the box. Flier points indicate outliers. The red horizontal dashed line indicates the experimental value for the percentage of coexisting pairs found in [6]. For the multi-species part the chosen parameters are $n=2000$, $N=100$, $R=0.2$, $N_{th}=60$, $K_{i}=1$ and $r_{i}=1\,\forall i$, $X^{0}_{i\alpha}=10^{-3}\,\forall\,(i,\,\alpha)$. The same for the two-species part except for $N=2$.
Figure 6: The simulations were conducted with $\sigma=0.1$ and varying $\mu$ across 0.1, 0.3, and 0.5 for the plots in the right column. Similarly, the simulations were repeated with $\mu=0.3$ while adjusting $\sigma$ to 0.05, 0.1, and 0.5 in the left column. The different colors of the circles represent the three values of $\mu$ in the right column and of $\sigma$ in the left column. The remaining parameters in the model were set as in Figure 2. The top row presents the Abundance Fluctuation Distribution, depicting the probability density for the rescaled log relative abundance of the species across 100 model runs. Here, the black line represents a gamma distribution. In the second row, we plotted the variance of the relative abundance as a function of the corresponding means across 100 simulations. The black line corresponds to Taylor’s law with an exponent of two. The third row displays the distribution for the rescaled log average relative abundance, again across 100 different simulations, compared with a lognormal distribution (black line). The analysis yielded mean $R^{2}$ values of 0.97, 0.5, and 0.86, respectively, for the three laws.
In fact, in the absence of multiple species, spatial self-organization in this scenario leads to a random process distributing both species equally across all patches with high probability. This implies an unaffected interaction pattern, rendering the spatial distribution incapable of promoting coexistence, in contrast to what happen in the case of multiple species, where the spatial simulation naturally generates more heterogeneous patches. When there is a significant disparity in the initial concentration of the two species, as previously mentioned, the patch structure does not impact the invasion criterion when one species invades the other. Consequently, the mesoscale structure, while potentially influencing coexistence in multi-species configurations, does not change the criteria for coexistence in the case of two species. This elucidates the findings illustrated in Figure 5.